OpenStudy (anonymous):

Find the absolute (global) Maximum and absolute (global) Minimum values of the function f(x)= (x^2)/(x^2-9) on the interval [-2,1].

OpenStudy (zarkon):

start by finding the critical numbers of \(f\)

OpenStudy (anonymous):

(-3,3)?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

there is only one critical number

OpenStudy (anonymous):

oh! it's 0

OpenStudy (zarkon):

correct

OpenStudy (zarkon):

and 0 falls between -2 and 1 so we keep it

OpenStudy (zarkon):

now plug -2,0, and 1 into \(f\) which is the biggest and which is the smallest

OpenStudy (anonymous):

f(-2)= -4/5 f(0)=0 f(1)=-1/8